Show that the closed sphere with centre )7,3,2( and radius 10 in 3 R is contained in the
open cube P = {(x, y,z :) x − 2 <11, y − 3 <11, z − 7 <11}.
1
Expert's answer
2020-02-25T11:25:30-0500

The equation of the closed sphere with the center "(2, 3, 7)" and radius "10" is


"(x-2)^2+(y-3)^2+(z-7)^2=10^2""| x-2|\leq 10, |y-3| \leq 10, |z-7|\leq 10""--10\leq x-2\leq 10, -10\leq y-3\leq 10, -10\leq z-7\leq 10"

Therefore the closed sphere with the center "(2,3,7)" and radius "10" is contained in the

open cube "P=\{(x,y,z):x-2<11,y-3<11,z-7<11\}".